OpenStudy (anonymous):

y varies directly as the square of x. When x = 1, y = 8. Find x when y = 128. x = ±4 x = ±6 x = ±8 x = ±16 what does this mean

OpenStudy (anonymous):

this means \[y=k \times x^2\]

OpenStudy (anonymous):

so\[8=k \times 1^2\]

OpenStudy (anonymous):

can you work out what the value of k is

OpenStudy (anonymous):

\[128=8\times x^2\] \[128\div8=x^2\] \[16=x^2\] x=4

OpenStudy (anonymous):

actually its \[\pm4\] as -4 squared = 16